OpenStudy (crashonce):

If the volume of a sphere is increased by 25%, its area is increased by how much?

OpenStudy (crashonce):

@ganeshie8

OpenStudy (crashonce):

@mathmath333

OpenStudy (crashonce):

@ParthKohli

OpenStudy (crashonce):

@Jhannybean

OpenStudy (crashonce):

@ganeshie8

OpenStudy (zzr0ck3r):

\(V=\frac{1}{3}rA\iff A=\frac{3V}r{}\)

OpenStudy (crashonce):

yes then

OpenStudy (zzr0ck3r):

?

OpenStudy (crashonce):

so what is the increase, can u provide more steps

OpenStudy (jhannybean):

\[A = 4\pi r^2\]\[V= \frac{4}{3} \pi r^3\]and radius = r \[A \propto r^2 ~,~ SA \propto r^3\]So if the volume increases by 25% (0.25), then the radius increases by \(\sqrt[3]{1.25}\) then our Area has increased by \((\sqrt[3]{1.25})^2\)

OpenStudy (mathmath333):

\(V=\dfrac{4}{3}\pi(r)^3\) let \(r=21\),then \(V=38808\) and \(V_1=1.25V\\ V_1=\dfrac{4}{3}\pi(r_1)^3 =48510\\ r_1=22.65\\ A=4\pi r^2\\ =4\pi (21)^2\\ =5544\\ A_1=4\pi (r_1)^2\\ =4\pi (22.65)^2\\ =6446.83\\ \%~increase=\dfrac{6446.83-5544}{5544}\times 100\\ \approx 16.28\% \)

OpenStudy (jhannybean):

sorry, typo.

OpenStudy (jhannybean):

\[r=\sqrt[3]{r}\]

OpenStudy (jhannybean):

r is radius..

OpenStudy (jhannybean):

@mathmath333 where did you get \(r=21\)?

OpenStudy (crashonce):

@Jhannybean what was ur solution

OpenStudy (mathmath333):

i assumed it to make the calulation with \(\pi\) easier

OpenStudy (jhannybean):

@CrashOnce calculate my answer and compare it with yours. I've set it up for you.

OpenStudy (jhannybean):

@mathmath333 we are off just a little bit, around the same answer though :)

OpenStudy (crashonce):

ill go with 16% thanks both

OpenStudy (jhannybean):

@CrashOnce , to find the percentage increase...

OpenStudy (jhannybean):

you take \(\left(\sqrt[3]{1.25}\right)^2\),and plug it back into your formula for Surface Area,

OpenStudy (mathmath333):

assumption method is useful when its objective type questions or the time is very less to solve the problem

OpenStudy (jhannybean):

\[A=4\pi r^2\]\[A=4\pi \left(\left(\sqrt[3]{1.25}\right)^2\right)^2\]

OpenStudy (jhannybean):

This will give you the percentage increase of Area.

OpenStudy (jhannybean):

Do you understand @CrashOnce ?

OpenStudy (crashonce):

yeas i do

OpenStudy (jhannybean):

Awesome :)